OpenStudy (unklerhaukus):

Macros \[\newcommand\dd[1]{\,\mathrm d#1} % infinitesimal \newcommand\de[2]{\frac{\mathrm d #1}{\mathrm d#2}} % first order derivative \newcommand\pa[2]{\frac{\partial #1}{\partial #2}} % partial derivative \newcommand\den[3]{\frac{\mathrm d^#3 #1}{\mathrm d#2^#3}} % n-th derivative \newcommand\pan[3]{\frac{\partial^#3 #1}{\partial#2^#3}} % n-th partial derivative \dd x \de yx \pa yx \den yxn \pan yxn\]

OpenStudy (unklerhaukus):

``` \newcommand\Beta[2]{\operatorname B \left(#1,#2\right)} % Beta function of (m,n) \Beta mn ``` \[\newcommand\Beta[2]{\operatorname B \left(#1,#2\right)} % Beta function of (m,n) \Beta mn\]

OpenStudy (unklerhaukus):

``` \newcommand\intl[4]{\int\limits_{#1}^{#2}{#3}{\dd#4}} % integral _{a}^{b}{f(x)}\dd x \newcommand\erf[1]{\operatorname{erf}\left(#1\right)} % Error function (#1) \newcommand\erfi[1]{\frac2{\sqrt\pi}\intl{0}{#1}{e^{-u^2}}{u}} % Error function integral (#1) \intl abfx\\ \erf x=\erfi x\\ \erf y=\erfi y\\ \erf z=\erfi z ``` \[\newcommand\intl[4]{\int\limits_{#1}^{#2}{#3}{\dd#4}} % integral _{a}^{b}{f(x)}\dd x \newcommand\erf[1]{\operatorname{erf}\left(#1\right)} % Error function (#1) \newcommand\erfi[1]{\frac2{\sqrt\pi}\intl{0}{#1}{e^{-u^2}}{u}} % Error function integral (#1) \intl abfx\\ \erf x=\erfi x\\ \erf y=\erfi y\\ \erf z=\erfi z\]